Grade 12Engineering Entrance Exams

An aqueous solution containing 4% nonvolatile solute exerts a pressure of 1.002 bar at the normal boiling point of the solvent. What is the molar mass of the solute?

sudhanshu
12 Years agoGrade 12

1 Answer

Avinash
11 Years ago

Vapour pressure of the solution at normal boiling point (p1) = 1.004 bar
Vapour pressure of pure water at normal boiling point
Mass of solute, (w2) = 2 g
Mass of solvent (water), (w1) = 98 g
Molar mass of solvent (water), (M1) = 18 g mol−1
According to Raoult’s law,
pi0-p/pi0=w2*m1/(w1*m2)
1.013-1.004/1.013=2*18/98*m2
= 41.35 g mol−1
Hence, the molar mass of the solute is 41.35 g mol−1.

question mark
Hello,
 
Can we learn and prepare for MHT-CET online ?
engineering entrance exams

Last Activity: 4 Years ago